OpenStudy (anonymous):

A triangle has side lengths of 34 in., 20 in., and 47 in. Classify it as acute, obtuse, or right.

hero (hero):

Use law of sines and cosines to find the angles

OpenStudy (bahrom7893):

i was thinking just draw it out or smth haha

hero (hero):

yup you could do that too using graph paper.

OpenStudy (bahrom7893):

not like 34 in, 34 millimeters maybe.

hero (hero):

I think that might be a little difficult though especially with having no clue about the angles

OpenStudy (bahrom7893):

It's def not right (just checked pythagorean thm)

hero (hero):

I bet @Mimi_x3 already knows the answer.

OpenStudy (mimi_x3):

oh gawd..again..i forgot how to do this. xD

hero (hero):

lol

OpenStudy (mimi_x3):

laughing at me? :(

hero (hero):

with you...

Directrix (directrix):

If the square of the longest side of a triangle is greater than the sum of the squares of the other two sides, then the triangle is obtuse. 47^2 = 2209; 34^2 = 1156; and 20^2 = 400. 2209 > 1156 + 400 2209 > 1556  The triangle is obtuse.

hero (hero):

I knew about that rule and didn't even think of it.